OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation.

OpenStudy (anonymous):

2x + 5y – 2z = –1 x + 3y + z = –2 3x – 2y + 3z = –17

OpenStudy (anonymous):

choices are: –y – 4z = 3 –11y = –11 11y = 3 7y + 6z = –11 4x + 11y = –5 –11y = –11 –y = –5 6x + 7y = –11

OpenStudy (anonymous):

I think it's the fourth choice but I'm not sure, can somebody please help me?